OpenStudy (anonymous):

Please please please help State whether the given measurements determine zero, one, or two triangles. A = 58°, a = 25, b = 28

OpenStudy (anonymous):



OpenStudy (anonymous):

You need to find at least another angle.

OpenStudy (anonymous):

You could use law of sines.

OpenStudy (anonymous):

B=71.77100453...?

OpenStudy (anonymous):

@Brittni0605 is that right? Thanks for your help so far

OpenStudy (anonymous):

I'd like to see your work.

OpenStudy (anonymous):

25/sin58=28/sinb 25/sin58 28/ANS sin-1(ANS) = 71.78

OpenStudy (anonymous):

Ok, yes.

OpenStudy (anonymous):

Thanks. So how do I use that to determine whether the given measurements determine zero, one, or two triangles

OpenStudy (anonymous):

What is the supplement angle to 71.78 degrees?

OpenStudy (anonymous):

I have no idea how to find that

OpenStudy (anonymous):

180-71.78 = ?

OpenStudy (anonymous):

oh. 108.22

OpenStudy (anonymous):

Good, so: 

OpenStudy (anonymous):

ok so its like the ambiguous case?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Since you know, one angle is 58. You have to check if the two possibilities for angle B are reasonable.

OpenStudy (anonymous):

so i do 25/sin58=28/sin71.78 and again with 25/sin58=28/sin108.22?

OpenStudy (anonymous):

No, you know all the angles in a triangle add up to 180, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So that means angle a, b, and c should sum to 180. you know angle a is 58. you have two possibilities for angle b, and well you dont know angle c.

OpenStudy (anonymous):

but since you know two angles, they cannot add to 180 bc you have another angle in the triangle. So to check if the two numbers for angle b are possible, you must..?

OpenStudy (anonymous):

add them?

OpenStudy (anonymous):

so only one is possible?

OpenStudy (anonymous):

because they equal 180?

OpenStudy (anonymous):

180 - (71.78+58) = .. 180 - (108.22 +58) = ..

OpenStudy (anonymous):

50.22 and 13.78

OpenStudy (anonymous):

Ok, so both are possible. Angle c could be 50.22 and 13.78..

OpenStudy (anonymous):

oh. so two are possible?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thanks. I think I actually get this a little better now. Thanks a ton.

OpenStudy (anonymous):

No prob.